OpenStudy (fixxer):

Help me find a vector parametric of the 2 intersecting surfaces. z=9-x^2-(y-1)^2 z=2y+4 See inside for more info

OpenStudy (fixxer):

I have tried setting z=z \[9-x^2-(y-1)^2-2y+4=x^2+y^2-4\] I then tried the following \[x=2\cos(t),y=2\sin(t)\]then i solved both surface equation substituting in for x and y\[z=4\sin(t)+4\] However this is not right, the spacecurve I get is similar in shape, but the orientation is wrong

OpenStudy (holsteremission):

Really? Looks fine to me. Or do you mean orientation as in the orientation of the path (clockwise vs counerclockwise)?

OpenStudy (fixxer):

that is odd I get something completely different :s

OpenStudy (fixxer):

Wow, I had really messed up the axes for the solid. Im glad you were here to confirm the parameters, after all i picked up a few of your trick :)

OpenStudy (holsteremission):

I think you might be using \(z=4-4\cos t\)?

OpenStudy (fixxer):

The problem was that I plotted y values on x axis and x values on y axsis for the solid. Didnt even slip my mind before you confirmed it was the right parameters. thanks

OpenStudy (holsteremission):

yw